# Quantum Statistical Effects on Warm Dark Matter and the Mass Constraint from the Cosmic Large Scale Structure

### v1

@@ 3 The mass constraints
It is now ready to investigate the effects of quantum statistics on the WDM mass constraint from the LSS. So far, the suppression on the small-scale matter power spectrum has not been observed and it thus puts a lower limit on the WDM mass. Recently, a strong constraint of $\ell_{\rm s}$ is given in Iršič et al. (2024) that the matter power spectrum cannot drop more than $5\%$ at the wavenumber of $14.35\,h$Mpc-1. This corresponds to $\ell_{\rm s}<0.135\,h^{-1}$pc. [^4] We have used $\left(\frac{P_{\rm WDM}}{P_{\rm CDM}}\right)^{1/2}\simeq\left(1+(\alpha k)^{2\mu}\right)^{-5/\mu}$ (Bode et al., 2001) with $\mu=1.12$ (Viel et al., 2005). We take $h=0.674$ and $\Omega_{\rm dm}h^{2}=0.12$ (Ade & others (2020), Planck Collaboration). According to Eq. (20), the upper bound of $\ell_{\rm s}$ is then translated into a lower bound of the WDM mass which depends on the degree of quantum statistics. In Figure 3 we show the constraints of the qWDM mass as a function of $\Delta_{\rm i}$ for the fermion case and $r_{\rm i}$ for the boson case.
@@
For the fiducial case, the mass is constrained to be $m>4.0$ keV. This constraint is somewhat different from $m>5.7$ keV given in Iršič et al. (2024) because we assume WDM keeps an equilibrium phase-space distribution instead of being a free stream; also see Egana-Ugrinovic et al. (2021). The classical case with $\Delta_{\rm i}\ll-1$ makes the mass lower bound stronger (i.e., higher). This is because the DM particle number is smaller in the classical case than in the fiducial case and then the $\beta$ factor is smaller. Meanwhile, the $C_{\rm cs}$ factor is similar as long as $\Delta_{\rm i}<0$. These result in a smaller $\gamma$ factor in Eq. (20) and a higher mass bound in the classical case compared to the fiducial case.
@@
For the fermion case with the presence of degenerate pressure, the mass constraint is only slightly lowered than the fiducial case. Two effects occur that almost cancel out each other and leave the mass constraint almost unchanged. First, DM is denser compared to the fiducial case, which raises the $\beta$ factor in Eq. (8) and tends to lower the mass constraint. But, as discussed in Sec. 2.2, RNRT is postponed due to the degenerate pressure so that $C_{\rm cs}$ is also larger. Both $\beta$ and $C_{\rm cs}^{3}$ have the same asymptotic dependence on $\Delta_{\rm i}$, i.e., $\propto\Delta^{3}_{\rm i}$ when $\Delta_{\rm i}\gg 1$. As a result, the $\gamma$ factor becomes a constant and the qWDM mass lower bound becomes independent on $\Delta_{\rm i}$ at the highly degenerate limit. Despite the different analyses here, the conclusion that fermionic degenerate pressure only slightly lowers the WDM mass constraint is consistent with other works; see (Bar et al., 2021; Carena et al., 2022).
@@
For the boson case with a BEC component, the mass constraint can be significantly lowered with a high BEC fraction. Similar to the fermion case, DM is denser compared to the fiducial case. However, different from the fermion case, now the transition is advanced. The net effect results in a larger $\gamma$ factor and a lowered mass bound. At a high BEC fraction, the WDM mass bound can be arbitrarily low. This clearly shows that DM constituted of condensation, such as that of the fuzzy DM (Ji & Sin, 1994; Hu et al., 2000), can avoid the WDM mass bound. However, we can see from the lower panel of Figure 3 that a high initial BEC fraction ($r_{\rm i}\gtrsim 0.9$) is needed to lower the mass bound significantly.
@@
Eventually at the low mass region for the boson case with a very high or even a 100% BEC component, the WDM mass will be restricted due to the “quantum pressure” (Chavanis, 2012; Iršič et al., 2017; Dalal & Kravtsov, 2022), the study of which is out of the scope of this work.
@@ 4 Remarks on quantum statistical effects during the transition
For the fermion case, $\Delta$ drops from one constant $\Delta_{\rm i}$ in the relativistic phase to another constant $\Delta_{\rm f}$ in the nonrelativistic phase. Thus, the degenerate pressure can only appear today if it appeared in the early universe. The drop of $\Delta$ is a result of the conservation of specific entropy. In both the relativistic and nonrelativistic phases, the specific entropy $s/n$ is a monotonic decreasing function only of $\Delta$. On the other hand, for a given $\Delta$, $s/n$ is smaller in the nonrelativistic phase than in the relativistic phase. Thus, to keep the specific entropy constant during RNRT, $\Delta$ must decrease. By equaling the relativistic and relativistic limits of $s/n$ [see Eq. (A9)], it can be shown that in the highly degenerate limit (i.e., $\Delta\rightarrow\infty$), $\Delta$ drops to half of its initial value, i.e., $\Delta_{\rm f}=\frac{1}{2}\Delta_{\rm i}$. So, if qWDM was initially in a highly degenerate state, it remains a highly degenerate state today.
@@
When $r_{\rm i}<r_{\rm i}^{\rm cr}$, the BEC fraction drops to 00 at some point during RNRT, and then $\Delta$ drops from 00 to a negative value until the nonrelativistic phase. On the other hand, if $r_{\rm i}\rightarrow 1$, we also have $r_{\rm f}\rightarrow 1$ from Eq. (22). Therefore, if qWDM started with a highly condensed state, it remains a highly condensed state today.
@@ 5 Summary and Conclusion
In this work, we have studied the quantum statistical effects of a single-species WDM on the early-time background evolution of the density and pressure, cosmic structure growth, and the mass constraint from the Lyman-$\alpha$ forest. We have considered situations where the system is degenerate for a fermion DM or where a BEC component appears for a boson DM. We have paid special attention to the change of the quantum statistical effects during RNRT, which affects the WDM sound horizon and the suppression scale of the matter power spectrum.
@@
When quantum statistical effects are considered, two factors predominantly affect the constraint on the WDM mass from LSS. One is the qWDM number density compared to the thermal relic WDM case. The other is the time of the qWDM relativistic-to-nonrelativistic transition. Situations with a higher qWDM number density and an earlier transition make the mass lower bound weaker (i.e., the bound is lower). In both the degenerate fermion case and the boson case with a BEC component, the number density is higher than in the thermal relic WDM case. For the fermion case, the degenerate pressure postpones RNRT. The net effect is the DM mass constraint is only slightly lower in the degenerate fermion case compared to that in the thermal relic WDM case. On the other hand, the boson case with a BEC component advances RNRT. As a result, the DM mass lower bound can be significantly relaxed. However, a large initial BEC fraction ($n_{\textsc{bec}}/n_{\rm tot}\gtrsim 90\%$) is required to substantially weaken the mass lower bound.
@@
The degree of degeneracy for the fermion case and the BEC fraction for the boson case decrease during RNRT, both of which are results of the conservation of specific entropy. In particular, for the boson case, the BEC fraction drops to zero unless the initial BEC fraction is higher than about 64%. Small-scale problems have motivated proposals that involve boson DM with a completely or partially condensed state, but dynamically generating a BEC component does not achieve such a high initial fraction (Madsen, 1992). This does not constitute immediate trouble to scenarios where the boson DM somehow directly starts with a completely condensed state as is what is taken for, e.g., axion DM (Preskill et al., 1983; Abbott & Sikivie, 1983; Dine & Fischler, 1983), Fuzzy DM (Ji & Sin, 1994; Hu et al., 2000) and others (Fukuyama & Morikawa, 2006). However, it poses a question on how to generate a high DM BEC fraction dynamically. To avoid the above constraint, one way is the BEC state is generated during the nonrelativistic phase of DM. Another way is that small-scale BEC can be somehow formed during gravitational collapse. A caveat in this work is that we assume qWDM is an ideal gas in an equilibrium phase-space distribution so that the background evolution can be established in a parameterized way. This implicitly puts a constraint on the type and the strength of the self-interaction. Considering more general self-interaction may result in different evolution (Harko, 2011; Guth et al., 2015).
@@
W. L. thanks Vid Iršič for answering the questions about the unit of wavenumber used in the Lyman-$\alpha$ forest flux power spectrum, Xingang Chen and Amol Upadh for useful feedback. W. L. acknowledges that this work is supported by the “Science & Technology Champion Project” (202005AB160002) and the “Top Team Project” (202305AT350002), both funded by the “Yunnan Revitalization Talent Support Program.” This work is also supported by the “Yunnan General Grant” (202401AT070489).
@@ Appendix A Methods
A.1 Thermodynamical variables and their relativistic and the nonrelativistic limits
@@
By defining the mass-to-temperature ratio $x\equiv m/T$ and the effective chemical potential-to-temperature ratio $\Delta\equiv(\mu-m)/T$, the above can be cast into following forms,
@@
where $n_{*}=\frac{gm^{3}}{2\pi^{2}\hbar^{3}}$. The relativistic and nonrelativistic limits of Eqs. (A4)-(A6) can be calculated analytically which we give in Table 1. Note that the condition for the nonrelativistic limit is $x\gg 1$ and $x\gg\Delta$ as we explained.
@@
Table 1: Summary of the relativistic and the non-relativistic limits of the characteristic functions of thermal number density ($n=\frac{gm^{3}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{3}}\mathcal{N}$), energy density ($\rho=\frac{gm^{4}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{4}}\mathcal{R}$), pressure ($p=\frac{gm^{4}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{4}}\mathcal{P}$), and specific entropy $s/n$ ($n$ is the thermal number density only). The relativistic (non-relativistic) limit is defined as $x\ll 1$ ($x\gg 1$ and $x\gg\Delta$). The function $f^{\mp}_{s}$ is defined as $f^{\mp}_{s}(\Delta)\equiv\sum_{n=0}^{\infty}\frac{(\mp 1)^{n}}{(n+1)^{s}}\exp(n\Delta)=\Phi(e^{\mp\Delta},s,1)$, where now -- is for fermion, $+$ for boson, and $\Phi$ is the Lerch transcendent. The functions ${\rm erfi}$ and $\rm{erfc}$ are the complementary and imaginary error functions. For the boson case with a BEC component, the above equations apply to the thermal component with $\Delta=0$. The integration forms of the $J$’s functions are defined as $J_{s}^{\mp}(\Delta)\equiv\int_{0}^{\infty}\frac{z^{s-1}{\rm d}z}{\exp(z-\Delta)\pm 1}$. Some important properties are $\frac{{\rm d}}{{\rm d}\Delta}J_{s}^{\mp}(\Delta)=(s-1)J_{s-1}^{\mp}(\Delta)$. Note that the subscript $s$ is an integer or half-integer not to be confused with the entropy density.
From Euler’s theorem on homogeneous functions, the entropy of an ideal gas is given by $TS=(U+VP-\mu N)$ for given volume $V$, energy $U$ (including the rest-energy), and particle number $N$. Dividing it by $V$ and $T$, we obtain the entropy density which reads,
@@
Note that while Eq. (A8) still holds for the boson case with a BEC component with $n$ being the density of the thermal component, the correct specific entropy that conserves is $s/n_{\rm tot}=(1-r)s/n$. We show the dependence of $s/n$ on $x$ and $\Delta$ in Figure 4. For a given $\Delta$, the value of $n/s$ is a constant in the relativistic limit and is a smaller constant in the nonrelativisitic limit. The large $|\Delta|$ behaviors of the relativistic and nonrelativistic limits of $s/n$ are useful and given by (positive $\Delta$ only applies to the fermion case),
@@
Figure 4: The dependence of thermal specific entropy $s/n$ on the mass-to-temperature ratio $x$ and the effective chemical potential-to-temperature ratio $\Delta$. Left: curves of $s/n$ as a function of $x$ for some fixed $\Delta$’s for the fermion case. For a given $\Delta$, $s/n$ drops from a constant in the relativistic limit to another constant in the nonrelativistic limit. Right: curves of $s/n$ as a function of $\Delta$ in the relativistic limit (solid) and in the relativistic limit (dashed). The blue curves are for fermion and the red curves are for boson. Note that, for the boson case, the specific entropy concerning the total particle number is $s/n_{\rm tot}=(1-r)s/n.$

### v2

@@ 3 The mass constraints
With the above discussions, it is now ready to see the effects of quantum statistics on the WDM mass constraint from the Lyman-$\alpha$ forest. So far, the suppression on the small-scale matter power spectrum has not been observed, and it thus puts a lower limit on the WDM mass. To apply the current Lyman-$\alpha$ forest constraints on the suppression scale, we first note that the “transfer function” of qWDM can be parameterized similarly to that of thermal relic WDM by (Bode et al., 2001)
$\left(\frac{P_{\rm WDM}}{P_{\rm CDM}}\right)^{\nicefrac{{1}}{{2}}}\simeq\left(1+(\alpha k)^{2\mu}\right)^{\nicefrac{{-5}}{{\mu}}}\,,$ (23)
with $\mu=1.12$ (Viel et al., 2005). With our definition of the suppression scale, we can relate $\alpha$ to $\ell_{\rm s}$ by
$\alpha=\frac{\ell_{\rm s}}{2\pi}\big{(}2^{\nicefrac{{\mu}}{{10}}}-1\big{)}^{\nicefrac{{1}}{{2\mu}}}\,.$ (24)
With $\ell_{\rm s}$ solved in Eq. (21), the parameterized transfer functions (squared) for different cases are represented by the dotted curves in Figure 3. The parameterized transfer function matches the overall suppression effect of the numerical result very well.
Recently, a strong constraint of $\ell_{\rm s}$ was given in Iršič et al. (2024) that the matter power spectrum cannot drop more than $5\%$ at the wavenumber of $14.35\,h$Mpc-1. This corresponds to $\ell_{\rm s}<0.135\,h^{-1}$pc by using Eq. (23). We take $h=0.674$ and $\Omega_{\rm dm}h^{2}=0.12$ (Ade & others (2020), Planck Collaboration). According to Eq. (21), the upper bound of $\ell_{\rm s}$ is then translated into a lower bound of the WDM mass that depends on the degree of quantum statistics. In Figure 4, we show the constraints of the qWDM mass as a function of $\Delta_{\rm i}$ for the fermion case and $r_{\rm i}$ for the boson case.
@@
For the fiducial case, the mass is constrained to be $m>4.6$ keV. This constraint is somewhat different from $m>5.7$ keV given in Iršič et al. (2024), as we assume WDM maintains an equilibrium phase-space distribution rather than being a free stream; also see Egana-Ugrinovic et al. (2021); Garani et al. (2022). Additionally, the precision of the usual mass-to-suppression scale parameterization at the large-mass region contributes to this difference (Vogel & Abazajian, 2023). The classical case with $\Delta_{\rm i}\ll-1$ makes the mass lower bound stronger (i.e., higher). This is because the DM particle number is smaller in the classical case than in the fiducial case, and then the $\beta$ factor is smaller. Meanwhile, the $C_{\rm cs}$ factor is similar as long as $\Delta_{\rm i}<0$. These result in a smaller $\gamma$ factor in Eq. (21) and a higher mass bound in the classical case compared to the fiducial case.
@@
For the fermion case with the presence of degenerate pressure, the mass constraint is only slightly lower than the fiducial case. Two effects occur that almost cancel each other out and leave the mass constraint almost unchanged. First, DM has a higher number density compared to the fiducial case for a given initial warmness, which raises the $\beta$ factor in Eq. (8) and tends to lower the mass constraint. But, as discussed in Sec. 2.2, RNRT is postponed due to the degenerate pressure so that $C_{\rm cs}$ is also larger. Both $\beta$ and $C_{\rm cs}^{3}$ have the same asymptotic dependence on $\Delta_{\rm i}$, i.e., $\propto\Delta^{3}_{\rm i}$ when $\Delta_{\rm i}\gg 1$. As a result, the $\gamma$ factor becomes a constant, and the lower bound of the qWDM mass becomes independent of $\Delta_{\rm i}$ at the highly degenerate limit. Despite the different analyses here, the conclusion that fermionic degenerate pressure only slightly lowers the WDM mass constraint is consistent with other works; see (Bar et al., 2021; Carena et al., 2022).
@@
For the boson case with a BEC component, the mass constraint can be significantly lowered with a high BEC fraction. Similar to the fermion case, DM is denser compared to the fiducial case. However, unlike the fermion case, the transition is now advanced. The net effect results in a larger $\gamma$ factor and a lower mass bound. At a high BEC fraction, the WDM mass bound can be arbitrarily low. This clearly shows that DM constituted of condensation, such as that of the fuzzy DM (Ji & Sin, 1994; Hu et al., 2000), can have a mass that is orders of magnitude lower than the WDM mass bound for the standard thermal scenario. However, we can see from the lower panel of Figure 4 that a high initial BEC fraction ($r_{\rm i}\gtrsim 0.9$) is needed to lower the mass bound significantly.
@@
Eventually, at the low-mass region for the boson case with a very high or even a 100% BEC component, the qWDM mass will be restricted due to the “quantum pressure,” such as the case for fuzzy DM (Chavanis, 2012; Iršič et al., 2017; Dalal & Kravtsov, 2022), although the exact mass constraint is model-dependent (Amin & Mirbabayi, 2024; Liu et al., 2024). The study of this region is beyond the scope of this work.
As shown in Figure 3, the primary effect of quantum statistics on the transfer function is a change in the suppression wavenumber and consequently in $\ell_{\rm s}$. By aligning the transfer functions, we can establish a correspondence between the qWDM mass and the thermal relic WDM mass, which reads
$m=0.84\,m_{\textsc{wdm}}\left(\frac{2/g}{\gamma(\Delta_{\rm i},r_{\rm i})}\right)^{\nicefrac{{1}}{{4}}}\,.$ (25)
The factor of $0.84$ arises from the fact that the suppression scale in the fiducial case with $m=0.84\,m_{\textsc{wdm}}$ matches the suppression scale in the thermal WDM case with $m_{\textsc{wdm}}$.
@@ 4 Remarks on quantum statistical effects during the transition
For the fermion case, $\Delta$ drops from one constant $\Delta_{\rm i}$ in the relativistic phase to another constant $\Delta_{\rm f}$ in the nonrelativistic phase. Thus, the degenerate pressure can only appear today if it appeared in the early Universe. The drop of $\Delta$ is a result of the conservation of specific entropy. In both the relativistic and nonrelativistic phases, the specific entropy $s/n$ is a monotonic decreasing function only of $\Delta$. On the other hand, for a given $\Delta$, $s/n$ is smaller in the nonrelativistic phase than in the relativistic phase. Thus, to keep the specific entropy constant during RNRT, $\Delta$ must decrease. By equaling the relativistic and relativistic limits of $s/n$ [see Eq. (A9)], it can be shown that in the highly degenerate limit (i.e., $\Delta\rightarrow\infty$), $\Delta$ drops to half of its initial value, i.e., $\Delta_{\rm f}=\frac{1}{2}\Delta_{\rm i}$. So, if qWDM was initially in a highly degenerate state, then it remains in a highly degenerate state today.
@@
When $r_{\rm i}<r_{\rm i}^{\rm cr}$, the BEC fraction drops to 00 at some point during RNRT, and then $\Delta$ drops from 00 to a negative value until the nonrelativistic phase. On the other hand, if $r_{\rm i}\rightarrow 1$, we also have $r_{\rm f}\rightarrow 1$ from Eq. (26). Therefore, if qWDM started with a highly condensed state, it remains in a highly condensed state today.
@@ 5 Summary and Conclusion
In this work, we have studied the quantum statistical effects of a single-species WDM on the early-time background evolution of the density and pressure, cosmic structure growth, and the mass constraint from the Lyman-$\alpha$ forest. We have considered situations where the system is degenerate for a fermion DM or where a BEC component appears for a boson DM. We have paid special attention to the change of quantum statistical effects during RNRT, which affects the WDM sound horizon and the suppression scale of the matter power spectrum.
@@
When quantum statistical effects are considered, two factors predominantly affect the constraint on the WDM mass from the cosmic structure. One is the qWDM number density compared to the thermal relic WDM case. The other is the time of the qWDM RNRT. Situations with a higher qWDM number density and an earlier transition make the mass lower bound weaker (i.e., the bound is lower). In both the degenerate fermion case and the boson case with a BEC component, the number density is higher than in the thermal relic WDM case. For the fermion case, the degenerate pressure postpones RNRT. The net effect is that the DM mass constraint is only slightly lower in the degenerate fermion case compared to that in the thermal relic WDM case. On the other hand, the boson case with a BEC component advances RNRT. As a result, the DM mass lower bound can be significantly relaxed. However, a large initial BEC fraction ($n_{\textsc{bec}}/n_{\rm tot}\gtrsim 90\%$) is required to substantially weaken the mass lower bound.
@@
The degree of degeneracy for the fermion case and the BEC fraction for the boson case decrease during RNRT, both of which are results of the conservation of specific entropy. In particular, for the boson case, the BEC fraction drops to zero unless the initial BEC fraction is higher than about 64%. Small-scale problems have motivated proposals that involve boson DM with a completely or partially condensed state, but dynamically generating a BEC component does not achieve such a high initial fraction (Madsen, 1992). This does not constitute immediate trouble in scenarios where the boson DM somehow directly starts with a completely condensed state, as is what is taken for, e.g., axion DM (Preskill et al., 1983; Abbott & Sikivie, 1983; Dine & Fischler, 1983), fuzzy DM (Ji & Sin, 1994; Hu et al., 2000) and others (Fukuyama & Morikawa, 2006). However, it poses a question on how to generate a high DM BEC fraction dynamically. One way to avoid the above constraint is that the BEC state is generated during the nonrelativistic phase of DM. Another way is that small-scale BEC can be somehow formed during gravitational collapse.
A caveat in this work is that we assume that qWDM is an ideal gas in an equilibrium phase-space distribution, so that the background evolution can be established in a parameterized way. This implicitly puts a constraint on the type and strength of the self-interaction. Considering more general self-interaction may result in different evolution (Harko, 2011; Guth et al., 2015). Importantly, our main conclusions – specifically, the variation in mass constraints with the degree of quantum statistics and the evolution of quantum statistics during RNRT – remain qualitatively robust, provided that qWDM remains in equilibrium until the completion of RNRT. Moreover, the necessary self-interaction strength is significantly below the upper limits set by current astrophysical constraints; see Appendix B.
@@
W. L. thanks Vid Iršič for answering the questions about the unit of wavenumber used in the Lyman-$\alpha$ forest flux power spectrum, and Xingang Chen and Amol Upadh for useful feedback. W. L. acknowledges that this work is supported by the “Science & Technology Champion Project” (202005AB160002) and the “Top Team Project” (202305AT350002), both funded by the “Yunnan Revitalization Talent Support Program.” This work is also supported by the “Yunnan General Grant” (202401AT070489).
@@ Appendix A Methods
A.1 Thermodynamical variables and their relativistic and their nonrelativistic limits
@@
By defining the mass-to-temperature ratio $x\equiv m/T$ and the effective chemical potential-to-temperature ratio $\Delta\equiv(\mu-m)/T$, the above can be cast into following forms:
@@
where $n_{*}=\frac{gm^{3}}{2\pi^{2}\hbar^{3}}$. The relativistic and nonrelativistic limits of Eqs. (A4)-(A6) can be calculated analytically, which we give in Table 1. Note that the condition for the nonrelativistic limit is $x\gg 1$ and $x\gg\Delta$ as we explained.
@@
Table 1: Summary of the relativistic and the nonrelativistic limits of the characteristic functions of thermal number density ($n=\frac{gm^{3}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{3}}\mathcal{N}$), energy density ($\rho=\frac{gm^{4}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{4}}\mathcal{R}$), pressure ($p=\frac{gm^{4}}{2\pi^{2}\hbar^{3}}\frac{1}{x^{4}}\mathcal{P}$), and specific entropy $s/n$ ($n$ is the thermal number density only). The relativistic (nonrelativistic) limit is defined as $x\ll 1$ ($x\gg 1$ and $x\gg\Delta$). The function $f^{\mp}_{s}$ is defined as $f^{\mp}_{s}(\Delta)\equiv\sum_{n=0}^{\infty}\frac{(\mp 1)^{n}}{(n+1)^{s}}\exp(n\Delta)=\Phi(e^{\mp\Delta},s,1)$, where now -- is for fermion, $+$ for boson, and $\Phi$ is the Lerch transcendent. The functions ${\rm erfi}$ and $\rm{erfc}$ are the complementary and imaginary error functions. For the boson case with a BEC component, the above equations apply to the thermal component with $\Delta=0$. The integration forms of the $J$’s functions are defined as $J_{s}^{\mp}(\Delta)\equiv\int_{0}^{\infty}\frac{z^{s-1}{\rm d}z}{\exp(z-\Delta)\pm 1}$. Some important properties are $\frac{{\rm d}}{{\rm d}\Delta}J_{s}^{\mp}(\Delta)=(s-1)J_{s-1}^{\mp}(\Delta)$. Note that the subscript $s$ is an integer or half-integer not to be confused with the entropy density.
From Euler’s theorem on homogeneous functions, the entropy of an ideal gas is given by $TS=(U+VP-\mu N)$ for given volume $V$, energy $U$ (including the rest energy), and particle number $N$. Dividing it by $V$ and $T$, we obtain the entropy density, which reads
@@
Note that while Eq. (A8) still holds for the boson case with a BEC component, with $n$ being the density of the thermal component, the correct specific entropy that conserves is $s/n_{\rm tot}=(1-r)s/n$. We show the dependence of $s/n$ on $x$ and $\Delta$ in Figure 5. For a given $\Delta$, the value of $n/s$ is a constant in the relativistic limit and is a smaller constant in the nonrelativistic limit. The large $|\Delta|$ behaviors of the relativistic and nonrelativistic limits of $s/n$ are useful and given by (positive $\Delta$ only applies to the fermion case),
@@
Figure 5: The dependence of thermal specific entropy $s/n$ on the mass-to-temperature ratio $x$ and the effective chemical potential-to-temperature ratio $\Delta$. Left: curves of $s/n$ as a function of $x$ for some fixed $\Delta$s for the fermion case. For a given $\Delta$, $s/n$ drops from a constant in the relativistic limit to another constant in the nonrelativistic limit. Right: curves of $s/n$ as a function of $\Delta$ in the relativistic limit (solid) and in the nonrelativistic limit (dashed). The blue curves are for fermion and the red curves are for boson. Note that, for the boson case, the specific entropy concerning the total particle number is $s/n_{\rm tot}=(1-r)s/n.$